OpenStudy (anonymous):

Lim Sin(x)/x^2+3x x->0

OpenStudy (anonymous):

\[\lim_{x\to 0}\frac{\sin(x)}{x}=1\]

OpenStudy (anonymous):

so \[\lim_{x\to 0}\frac{\sin(x)}{x(x+3)}\] is what you have to compute break it in to two parts

OpenStudy (anonymous):

i guess to be more precise i should say "write as a product"

OpenStudy (anonymous):

so you get \[\lim_{x \rightarrow 0} \frac{ 1 }{(x+3) } =1 ??\]

OpenStudy (anonymous):

so the limit is 1/3?

OpenStudy (anonymous):

your answer is correct, but i assume you meant \[\lim_{x \rightarrow 0} \frac{ 1 }{(x+3) } =\frac{1}{3} \]

OpenStudy (anonymous):

yea ahah thanks i forgot the /3 thank you!

OpenStudy (anonymous):

yw